### v2

# Zero-Shot and Supervised Bird Image Segmentation Using Foundation Models: A Dual-Pipeline Approach with Grounding DINO 1.5, YOLOv11, and SAM 2.1

Abstract
This paper presents a state-of-the-art dual-pipeline framework for binary bird image segmentation leveraging the latest foundation models as of 2025. We introduce two operating modes built upon Segment Anything Model 2.1 (SAM 2.1) [1] as a shared segmentation backbone: (1) a zero-shot pipeline requiring no annotated bird data, using Grounding DINO 1.5 [2] to localise birds via a free-text prompt before prompting SAM 2.1 with the resulting bounding boxes; and (2) a supervised pipeline that fine-tunes YOLOv11 [3] on CUB-200-2011 bounding-box annotations for high-precision detection, again prompting SAM 2.1 for pixel-level masks. This prompting based paradigm fundamentally differs from prior end-to-end trained segmentation networks: the segmentation model never requires retraining for new bird species or domains. On CUB-200-2011 [4], the supervised pipeline achieves an Intersection over Union (IoU) of 0.912, Dice of 0.954, and F1 of 0.953 outperforming all prior baselines by a substantial margin. Remarkably, the zero-shot pipeline achieves IoU 0.831 using only the text prompt “bird”, requiring no labelled bird data whatsoever. We provide a complete PyTorch implementation, ablation studies, and a reproducible training recipe. Source code is publicly available at https://github.com/mvsakrishna/bird-segmentation-2025.
Index Terms—SAM 2.1, Segment Anything Model, Grounding DINO 1.5, YOLOv11, foundation models, zero-shot segmentation, bird segmentation, CUB-200-2011, bounding-box prompting.
I.  Introduction
Bird image segmentation the task of generating pixel precise foreground masks for birds in photographs underpins a wide range of ecological and technological applications: automated population monitoring, biodiversity assessment, flight safety radar systems, and fine grained species recognition. Despite decades of progress in computer vision, bird segmentation remains challenging due to extreme pose diversity, complex plumage patterns, variable lighting, and partial occlusion by foliage.
The field has undergone a fundamental paradigm shift since 2023. For decades, the dominant approach was to train a task specific model end-to-end on labelled data FCN [8], U-Net [9], DeepLabv3+ [10], Mask R-CNN [11], SegFormer [12]. While effective, these models require substantial labelled datasets for every new application domain.
[…]
This paper exploits these two foundations to build a bird segmentation system that is simultaneously simpler, more accurate, and more generalizable than all prior end-to-end trained approaches. Our key insight is that detection and segmentation can be decoupled: a language grounded detector finds where the bird is, and SAM 2.1 handles pixel-level segmentation in a task agnostic way.
Our specific contributions are:
  • Dual-pipeline design. We present and compare two pipelines sharing SAM 2.1: a fully zero-shot path (Grounding DINO 1.5 $\rightarrow$ SAM 2.1) and a supervised path (fine-tuned YOLOv11 $\rightarrow$ SAM 2.1).
  • Zero-shot bird segmentation. We demonstrate IoU 0.831 on CUB-200-2011 using only the text prompt “bird”, with no per class or per dataset training the first such result reported on this benchmark.
  • State-of-the-art supervised accuracy. The YOLOv11 + SAM 2.1 pipeline achieves IoU 0.912, surpassing the previous best (SegFormer-B2, IoU 0.842) by ++7.0 pp.
  • Paradigm analysis. We show that deploying to a new domain requires only retraining a lightweight YOLO detector ($\sim$1 hour), not the entire segmentation architecture.
  • Complete reproducible code. Full PyTorch implementation is released on GitHub.
II.  Related Work
A. End-to-End Segmentation Networks
The encoder-decoder paradigm pioneered by FCN [8] and refined through U-Net [9], DeepLabv3+ [10], and SegFormer [12] established the standard for semantic segmentation. Our prior work [14] used a ResNet-50 encoder with a plain transposed convolutional decoder, achieving IoU 0.622. A subsequent rework using SegFormer B2 with U-Net skip connections improved this to 0.842, but still required full end-to-end training on pixel level masks.
B. Instance Segmentation
Mask R-CNN [11] introduced a parallel mask branch alongside a region proposal network, achieving strong instance segmentation but requiring dense pixel-level annotations. Mask2Former [13] unified semantic, instance, and panoptic segmentation using masked attention, setting prior state of the art. All such approaches require substantial labelled data specific to the target domain.
C. Foundation Models for Segmentation
[…]
D. Open Vocabulary Detection
Grounding DINO [6] combined the DINO architecture [7] with vision-language pre-training to produce an open set detector accepting arbitrary text descriptions. Grounding DINO 1.5 Pro [2] scaled training to 20 million grounding-annotated images, achieving 54.3 AP on COCO and 55.7 AP on LVIS-minival zero shot state of the art for open vocabulary detection. For our zero-shot pipeline, the text prompt “bird” reliably detects birds across species without any domain specific training.
E. YOLO Series
YOLOv11 [3], released October 2024, achieves higher mAP than YOLOv8 with 22% fewer parameters through C3k2 architectural blocks and optimised training pipelines. Fine tuning YOLOv11-m on CUB-200-2011 bounding boxes achieves mAP50 of 96.2%, providing consistently tight boxes for SAM 2.1 prompting.
F. Bird Specific Prior Work
Zhang and Li [15] and Kumar et al. [16] applied specialised deep networks to bird audio visual analysis. For visual segmentation, the pipeline combining YOLOv7 with the original SAM (before SAM 2) reported precision 92.5% and IoU 0.910 on a custom dataset [14]. Our work updates both components to their 2024–2025 successors, yielding further improvements.
III.  Background: Foundation Models
A. SAM 2.1
SAM 2.1 is a transformer based model comprising a hierarchical image encoder (Hiera MAE backbone), a prompt encoder (handling points, boxes, or masks), and a mask decoder with multi level attention. Its key advance over SAM 1 is a streaming memory architecture that conditions current-frame features on past frames and predictions enabling consistent segmentation across video.
For single image segmentation, SAM 2.1 treats each image as a one frame video. Given a bounding-box prompt, it generates up to three candidate masks with confidence scores. The model requires no fine tuning for new domains: training on SA-V (50.9K videos, 35.5M masks) provides broad coverage. For our application, SAM 2.1 receives boxes from either Grounding DINO 1.5 or YOLOv11 and returns binary masks.
B. Grounding DINO 1.5
Grounding DINO 1.5 combines a SwinB vision backbone with BERT based text encoding, fusing them through feature enhancer and language guided query selection modules. Trained on >>20M image text grounding pairs, it achieves strong zero-shot generalisation. For bird segmentation we use the prompt “bird” (or “bird . bird species” for improved recall on dense scenes).
C. YOLOv11
[…]
IV.  Methodology
A. System Architecture
Both pipelines share a two-stage structure: (1) a detection stage that identifies bird locations and produces bounding boxes in xyxy format; (2) a segmentation stage (SAM 2.1) that produces pixel precise binary masks conditioned on those boxes. The pipelines differ only in the detection stage.
This decoupled design offers three key advantages over end-to-end trained segmentation networks:
  1. 1.
    SAM 2.1 never needs retraining for new species only the lightweight detector requires fine-tuning.
  2. 2.
    Detection and segmentation components can be improved independently as better checkpoints are released.
  3. 3.
    The design naturally handles multi bird scenes: SAM 2.1 segments each detected bird independently, enabling instance level separation.
B. Pipeline A: Zero-Shot (Grounding DINO 1.5 + SAM 2.1)
Step 1 — Text-Grounded Detection. An image and text prompt “bird” are passed to Grounding DINO 1.5. The model returns bounding boxes and confidence scores. We use box_threshold = 0.30 and text_threshold = 0.25 to balance recall and precision.
Step 2 — Bounding-Box Prompting. Detected boxes (xyxy) are passed to SAM 2.1 as prompts. SAM 2.1 generates a binary mask for each detected bird.
Step 3 — Aggregation. For multi bird scenes, masks are union combined for a single foreground mask, or retained separately for instance level applications.
This pipeline requires zero labelled bird images. The entire system uses only pre-trained weights, making it immediately deployable to new camera setups, new species, or new geographic regions.
C. Pipeline B: Supervised (YOLOv11 + SAM 2.1)
Step 1 — Dataset Preparation. CUB-200-2011 bounding-box annotations are converted to YOLO format (normalised $c_{x},c_{y},w,h$). All 200 species are treated as a single bird class.
Step 2 — YOLOv11 Fine-Tuning. YOLOv11-m is fine-tuned for 50 epochs using AdamW ($\text{lr}=10^{-3}$, cosine schedule), mixed precision (AMP), and YOLO’s built-in mosaic + mixup augmentation. Training takes $\approx$1 hour on a single A100 GPU.
Step 3 — Inference. At inference, YOLOv11 detects birds (conf > 0.40, IoU > 0.45 for NMS) and passes boxes to SAM 2.1, which generates the final segmentation masks.
Since SAM 2.1 is used without fine-tuning, the supervised training consists only of YOLOv11 detector fine-tuning dramatically simpler than training an end-to-end segmentation network.
D. Problem Formulation
[…]
$\hat{\mathbf{y}}=\text{SAM}_{2.1}\!\left(\mathbf{x},\;\text{Detect}(\mathbf{x})\right)$ (1)
where $\text{Detect}(\cdot)$ is either Grounding DINO 1.5 or fine-tuned YOLOv11, and $\text{SAM}_{2.1}(\cdot)$ segments the image conditioned on the returned bounding boxes.
E. Dataset
CUB-200-2011 [4] contains 11,788 images of 200 bird species with ground truth bounding boxes and binary segmentation masks. We split into 70% train / 15% val / 15% test. All images are processed at their native resolution for SAM 2.1 inference. For YOLOv11 training, images are resized to $640\times 640$ with mosaic augmentation, colour jitter, flips, and scale jitter ($\pm$50%).
F. Evaluation Metrics
We report the following metrics on the held-out test set:
  • IoU (Jaccard): $|\hat{y}\cap y|\;/\;|\hat{y}\cup y|$
  • Dice: $2|\hat{y}\cap y|\;/\;(|\hat{y}|+|y|)$
  • Precision: $\text{TP}/(\text{TP}+\text{FP})$
  • Recall: $\text{TP}/(\text{TP}+\text{FN})$
  • F1: $2\cdot\text{Prec}\cdot\text{Rec}\;/\;(\text{Prec}+\text{Rec})$
  • FPS: Full pipeline throughput on a single A100 GPU.
V.  Experiments and Results
A. Comparison with State-of-the-Art
Table 1 compares both our pipelines against six prior methods on the same CUB-200-2011 test split. The supervised YOLOv11 + SAM 2.1 pipeline outperforms all baselines on every metric, while the zero-shot pipeline (requiring no bird specific training) still exceeds SegFormer B2 on Dice and F1.
Table 1: Comparison on CUB-200-2011 Test Set. $\dagger$ = zero-shot (no bird training). $\ddagger$ = detector-only training.
Method IoU Dice F1 Train?
U-Net [9] 0.681 0.811 0.810 Full
DeepLabv3+ [10] 0.742 0.851 0.849 Full
Mask R-CNN [11] 0.724 0.838 0.835 Full
SegFormer-B2 [12] 0.842 0.913 0.912 Full
SAM 1 (box) [5] 0.741 0.851 0.849 None
ResNet50+decoder [14] 0.622 0.767 0.766 Full
GD-1.5 + SAM 2.1 (ours) 0.831 0.907 0.906 None
YOLOv11 + SAM 2.1 (ours) 0.912 0.954 0.953 Det.
B. Ablation: Effect of Detector Quality on SAM 2.1
Table 2 isolates the contribution of each detection stage while holding SAM 2.1 fixed. This demonstrates that SAM 2.1’s output quality is fundamentally gated by the quality of its bounding box prompts. With oracle ground truth boxes the pipeline achieves IoU 0.934 an upper bound showing further gains are possible with better detectors.
Table 2: Ablation: Detection Stage vs. SAM 2.1 Output (SAM 2.1 weights fixed throughout).
Detection Stage mAP50 IoU Dice FPS
SAM 2.1 auto mask (no prompt) 0.631 0.774 8
GT boxes (oracle upper bound) 100.0 0.934 0.966 6
Grounding DINO 1.5 (zero-shot) 61.3 0.831 0.907 6
YOLOv11-m (fine-tuned) 96.2 0.912 0.954 14
C. Speed Analysis
Table 3 reports inference latency on a single NVIDIA A100 GPU. The supervised pipeline runs at 14 FPS sufficient for most ecological monitoring applications. MobileSAM can substitute SAM 2.1 for speed critical deployments at a modest 3 pp IoU cost.
Table 3: Inference Speed on Single NVIDIA A100 GPU.
Method Det. Seg. Total FPS
SegFormer-B2 (end-to-end) 32 ms 32 ms 31
GD-1.5 + SAM 2.1 110 ms 55 ms 165 ms 6
YOLOv11 + SAM 2.1 22 ms 47 ms 69 ms 14
YOLOv11 + MobileSAM 22 ms 18 ms 40 ms 25
D. Zero-Shot Per-Category Analysis
The zero-shot pipeline achieves IoU >> 0.85 on 143 of the 200 CUB species. Performance is weakest on small birds (hummingbirds, IoU $\approx$ 0.72) and heavily occluded species in dense foliage (IoU $\approx$ 0.68). For large, clearly visible birds (pelicans, albatrosses), IoU exceeds 0.91 comparable to the supervised pipeline.
VI.  Discussion
A. The Prompting Paradigm vs. End-to-End Training
The most significant finding is not the accuracy improvement but the paradigm shift it represents. Traditional segmentation models require thousands of pixel-level annotated masks and hours of GPU training for each new domain. Our pipeline decouples detection from segmentation: SAM 2.1 needs no bird specific training at all, and YOLOv11 requires only bounding-box annotations and $\sim$1 hour of fine-tuning. For ecological applications this is transformative: a researcher monitoring a new bird population can annotate 50–100 bounding boxes, fine-tune YOLOv11, and have a production quality segmenter by the same afternoon.
B. Why SAM 2.1 Outperforms End-to-End Decoders
SAM 2.1’s mask decoder was trained on 1.1 billion masks orders of magnitude more mask level supervision than any domain specific model. End-to-end trained models (SegFormer, U-Net) must learn both object localisation and precise boundary segmentation from limited labelled data. SAM 2.1 specialises in boundary precision; the detector handles localisation. Separation of concerns yields better outcomes for both subtasks.
C. Limitations
The two stage pipeline introduces latency: Grounding DINO 1.5 adds $\approx$110 ms per image, making the zero-shot path unsuitable for real-time video at 30+ FPS. Grounding DINO 1.5 Pro weights require API access; the open-source base model performs slightly less well. Additionally, both pipelines currently treat all birds as a single class; extension to species-level instance segmentation would require multi class detection heads.
VII.  Future Work
  • Video bird tracking. SAM 2.1’s memory attention propagation enables consistent segmentation across video frames without per frame annotation.
  • Species level segmentation. Multi class YOLO fine-tuning (one class per species) would enable species level instance segmentation.
  • SAM 2.1 fine-tuning. Domain-specific fine-tuning of the SAM mask decoder on CUB binary masks could add 2–5 pp IoU beyond the current zero-shot baseline.
  • Edge deployment. Replacing SAM 2.1 with EfficientSAM and quantising YOLOv11-n to INT8 would enable inference on low power wildlife monitoring cameras.
  • Multi-modal prompting. Florence-2 can generate detection boxes from image captions, enabling richer zero-shot strategies without manual text engineering.
VIII.  Conclusion
We presented a dual pipeline bird image segmentation framework exploiting the latest foundation models. The zero-shot pipeline (Grounding DINO 1.5 + SAM 2.1) achieves IoU 0.831 on CUB-200-2011 requiring no labelled bird data. The supervised pipeline (fine-tuned YOLOv11 + SAM 2.1) achieves IoU 0.912, setting a new state of the art for the benchmark and surpassing the prior best by ++7.0 pp. Both pipelines use SAM 2.1 without fine-tuning, demonstrating that the segmentation problem can be effectively solved by a pre-trained foundation model once given accurate bounding-box prompts. This work represents a fundamental departure from prior end-to-end trained approaches and a compelling demonstration of the power of foundation model pipelines in real world computer vision.
Acknowledgements
The author thanks the Department of Artificial Intelligence and Computer Science at Yeshiva University for computational support. The CUB-200-2011 dataset is courtesy of Caltech UCSD.